# VERSA: A Versatile Evaluation Toolkit for Speech, Audio, and Music

### v1

@@ 5 Conclusion
Through its integration of more than 63 metrics and 711 variants, VERSA provides unparalleled support for evaluating speech synthesis, audio coding, music generation, and more. The toolkit not only simplifies the evaluation process but also addresses challenges such as bias in subjective evaluations and the need for domain-specific expertise.
@@ A.1 Base Configuration
1# Example YAML config
2
3# mcd f0 related metrics
4# -- mcd: mel cepstral distortion
5# -- f0_corr: f0 correlation
6# -- f0_rmse: f0 root mean square error
7- name: mcd_f0
8 f0min: 40
9 f0max: 800
10 mcep_shift: 5
11 mcep_fftl: 1024
12 mcep_dim: 39
13 mcep_alpha: 0.466
14 seq_mismatch_tolerance: 0.1
15 power_threshold: -20
16 dtw: false
17
18# signal related metrics
19# -- sir: signal to interference ratio
20# -- sar: signal to artifact ratio
21# -- sdr: signal to distortion ratio
22# -- ci-sdr: scale-invariant signal to distortion ratio
23# -- si-snri: scale-invariant signal to noise ratio improvement
24- name: signal_metric
25
26# pesq related metrics
27# -- pesq: perceptual evaluation of speech quality
28- name: pesq
29
30# stoi related metrics
31# -- stoi: short-time objective intelligibility
32- name: stoi
33
34# discrete speech metrics
35# -- speech_bert: speech bert score
36# -- speech_bleu: speech bleu score
37# -- speech_token_distance: speech token distance score
38- name: discrete_speech
39
40# pseudo subjective metrics
41# -- utmos: UT-MOS score
42# -- dnsmos: DNS-MOS score
43# -- plcmos: PLC-MOS score
44- name: pseudo_mos
45 predictor_types: ["utmos", "dnsmos", "plcmos", "singmos"]
46 predictor_args:
47 utmos:
48 fs: 16000
49 dnsmos:
50 fs: 16000
51 plcmos:
52 fs: 16000
53
54# speaker related metrics
55# -- spk_similarity: speaker cosine similarity
56# model tag can be any ESPnet-SPK huggingface repo at
57# https://huggingface.co/espnet
58- name: speaker
59 model_tag: default
60
61# torchaudio-squim
62# -- torch_squim_pesq: reference-less pesq
63# -- torch_squim_stoi: reference-less stoi
64# -- torch_squim_si_sdr: reference-less si-sdr
65# -- torch_squim_mos: MOS score with reference
66- name: squim_ref
67- name: squim_no_ref
68
69# An overall model on MOS-bench from Sheet toolkit
70# More info in https://github.com/unilight/sheet/tree/main
71# --sheet_ssqa: the mos prediction from sheet_ssqa
72- name: sheet_ssqa
73
74# Word error rate with OpenAI-Whisper model
75# More model_tag can be from https://github.com/openai/whisper?tab=readme-ov-file#available-models-and-languages .
76# The default model is ‘large-v3‘.
77# NOTE(jiatong): further aggregation are necessary for corpus-level WER/CER
78# --whisper_hyp_text: the hypothesis from ESPnet ASR decoding
79# --ref_text: reference text (after cleaner)
80# --whisper_wer_delete: delete errors
81# --whisper_wer_insert: insertion errors
82# --whisper_wer_replace: replacement errors
83# --whisper_wer_equal: correct matching words/character counts
84# --whisper_cer_delete: delete errors
85# --whisper_cer_insert: insertion errors
86# --whisper_cer_replace: replacement errors
87# --whisper_cer_equal: correct matching words/character counts
88- name: whisper_wer
89 model_tag: default
90 beam_size: 5
91 text_cleaner: whisper_basic
92
93# Speech Enhancement-based Metrics
94# model tag can be any ESPnet-SE huggingface repo
95# -- se_sdr: the SDR from a reference speech enhancement model
96# -- se_sar: the SAR from a reference speech enhancement model
97# -- se_ci_sdr: the CI-SDR from a reference speech enhancement model
98# -- se_si_snr: the SI-SNR from a rerference speech enhancement model
99- name: se_snr
100 model_tag: default
@@ Appendix B Additional System Design
Codec kbps SDR($\uparrow$) PESQ($\uparrow$) UTMOS($\uparrow$) DNSMOS (P.835)($\uparrow$) SPK-SIM($\uparrow$)
Encodec+ Défossez et al. (2024a) 6.00 6.88 2.77 3.09 2.96 0.72
DAC Kumar et al. (2024) 6.00 3.42 3.40 3.60 3.16 0.73
SpeechTokenizer Zhang et al. (2024b) 4.00 4.06 2.62 3.84 3.17 0.86
SQ-Codec Yang et al. (2024) 8.00 10.84 4.24 4.05 3.21 0.96
SoundStream (ESPnet) Shi et al. (2024) 4.00 -1.28 2.86 3.61 3.13 0.89
Speech-DAC (ESPnet) Shi et al. (2024) 4.00 6.35 3.10 3.87 3.23 0.87
Mimi+ Défossez et al. (2024b) 4.40 9.64 3.38 3.92 3.18 0.85
BigCodec Xin et al. (2024) 1.04 4.26 2.68 4.11 3.26 0.62
WavTokenizer Ji et al. (2024) 0.98 2.51 1.88 3.77 3.18 0.60
TS3-Codec+ Wu et al. (2024a) 0.85 2.94 2.23 3.84 3.21 0.70
Ground Truth - - - 4.09 3.18 -
@@
Community-driven Contribution: VERSA is designed not just as a toolkit, but as an evolving, community-driven platform that thrives on collaboration and shared innovation. To encourage and facilitate contributions, we provide clear and comprehensive guidelines for community members, available at https://github.com/shinjiwlab/versa/blob/main/contributing.md. These guidelines outline the process for proposing new features, adding metrics, reporting issues, and improving the existing codebase. By fostering an open and inclusive environment, VERSA invites researchers, developers, and practitioners from diverse backgrounds to contribute their expertise, ensuring that the toolkit remains up-to-date and relevant for a wide range of applications. This community-driven approach not only accelerates the development of VERSA but also ensures that it reflects the collective needs and insights of the global sound generation research community. Through active collaboration, VERSA aims to establish itself as a dynamic and enduring resource for the advancement of sound evaluation technologies.
@@ C.3 Speech Enhancement Evaluation Demonstration
The evaluation of speech enhancement is demonstrated in Table 7. While this paper focuses on six specific metrics, we anticipate broader adoption of the task by leveraging the diverse set of 63 metrics available in VERSA.

### v2

@@ 5 Conclusion
Through its integration of more than 65 metrics and 729 variants, VERSA provides unparalleled support for evaluating speech synthesis, audio coding, music generation, and more. The toolkit not only simplifies the evaluation process but also addresses challenges such as bias in subjective evaluations and the need for domain-specific expertise.
@@ 7.2 Future Adaptation
Acknowledgements
This work is supported by the Defence Science and Technology Agency (DSTA) in Singapore. We would like to thank Daniel Leong and Megan Choo for their valuable comments. Experiments of this work used the Bridges2 at PSC and Delta/DeltaAI NCSA computing systems through allocation CIS210014 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, supported by National Science Foundation grants 2138259, 2138286, 2138307, 2137603, and 2138296.
@@ A.1 Base Configuration
# Example YAML config
# mcd f0 related metrics
# -- mcd: mel cepstral distortion
# -- f0_corr: f0 correlation
# -- f0_rmse: f0 root mean square error
- name: mcd_f0
f0min: 40
f0max: 800
mcep_shift: 5
mcep_fftl: 1024
mcep_dim: 39
mcep_alpha: 0.466
seq_mismatch_tolerance: 0.1
power_threshold: -20
dtw: false
# signal related metrics
# -- sir: signal to interference ratio
# -- sar: signal to artifact ratio
# -- sdr: signal to distortion ratio
# -- ci-sdr: scale-invariant signal to distortion ratio
# -- si-snri: scale-invariant signal to noise ratio improvement
- name: signal_metric
# pesq related metrics
# -- pesq: perceptual evaluation of speech quality
- name: pesq
# stoi related metrics
# -- stoi: short-time objective intelligibility
- name: stoi
# discrete speech metrics
# -- speech_bert: speech bert score
# -- speech_bleu: speech bleu score
# -- speech_token_distance: speech token distance score
- name: discrete_speech
# pseudo subjective metrics
# -- utmos: UT-MOS score
# -- dnsmos: DNS-MOS score
# -- plcmos: PLC-MOS score
- name: pseudo_mos
predictor_types: ["utmos", "dnsmos", "plcmos", "singmos"]
predictor_args:
utmos:
fs: 16000
dnsmos:
fs: 16000
plcmos:
fs: 16000
# speaker related metrics
# -- spk_similarity: speaker cosine similarity
# model tag can be any ESPnet-SPK huggingface repo at
# https://huggingface.co/espnet
- name: speaker
model_tag: default
# torchaudio-squim
# -- torch_squim_pesq: reference-less pesq
# -- torch_squim_stoi: reference-less stoi
# -- torch_squim_si_sdr: reference-less si-sdr
# -- torch_squim_mos: MOS score with reference
- name: squim_ref
- name: squim_no_ref
# An overall model on MOS-bench from Sheet toolkit
# More info in https://github.com/unilight/sheet/tree/main
# --sheet_ssqa: the mos prediction from sheet_ssqa
- name: sheet_ssqa
# Word error rate with OpenAI-Whisper model
# More model_tag can be from https://github.com/openai/whisper?tab=readme-ov-file#available-models-and-languages .
# The default model is ‘large-v3‘.
# NOTE(jiatong): further aggregation are necessary for corpus-level WER/CER
# --whisper_hyp_text: the hypothesis from ESPnet ASR decoding
# --ref_text: reference text (after cleaner)
# --whisper_wer_delete: delete errors
# --whisper_wer_insert: insertion errors
# --whisper_wer_replace: replacement errors
# --whisper_wer_equal: correct matching words/character counts
# --whisper_cer_delete: delete errors
# --whisper_cer_insert: insertion errors
# --whisper_cer_replace: replacement errors
# --whisper_cer_equal: correct matching words/character counts
- name: whisper_wer
model_tag: default
beam_size: 5
text_cleaner: whisper_basic
# Speech Enhancement-based Metrics
# model tag can be any ESPnet-SE huggingface repo
# -- se_sdr: the SDR from a reference speech enhancement model
# -- se_sar: the SAR from a reference speech enhancement model
# -- se_ci_sdr: the CI-SDR from a reference speech enhancement model
# -- se_si_snr: the SI-SNR from a rerference speech enhancement model
- name: se_snr
model_tag: default
@@ Appendix B Additional System Design
Codec kbps PESQ($\uparrow$) UTMOS($\uparrow$) DNSMOS (P.835)($\uparrow$) SPK-SIM($\uparrow$)
Encodec+ Défossez et al. (2024a) 6.00 2.77 3.09 2.96 0.72
DAC Kumar et al. (2024) 6.00 3.40 3.60 3.16 0.73
SpeechTokenizer Zhang et al. (2024b) 4.00 2.62 3.84 3.17 0.86
SQ-Codec Yang et al. (2024) 8.00 4.24 4.05 3.21 0.96
SoundStream (ESPnet) Shi et al. (2024) 4.00 2.86 3.61 3.13 0.89
Speech-DAC (ESPnet) Shi et al. (2024) 4.00 3.10 3.87 3.23 0.87
Mimi+ Défossez et al. (2024b) 4.40 3.38 3.92 3.18 0.85
BigCodec Xin et al. (2024) 1.04 2.68 4.11 3.26 0.62
WavTokenizer Ji et al. (2024) 0.98 1.88 3.77 3.18 0.60
TS3-Codec+ Wu et al. (2024a) 0.85 2.23 3.84 3.21 0.70
Ground Truth - - 4.09 3.18 -
@@
Community-driven Contribution: VERSA is designed not just as a toolkit, but as an evolving, community-driven platform that thrives on collaboration and shared innovation. To encourage and facilitate contributions, we provide clear and comprehensive guidelines for community members, available at https://github.com/wavlab-speech/versa/blob/main/contributing.md. These guidelines outline the process for proposing new features, adding metrics, reporting issues, and improving the existing codebase. By fostering an open and inclusive environment, VERSA invites researchers, developers, and practitioners from diverse backgrounds to contribute their expertise, ensuring that the toolkit remains up-to-date and relevant for a wide range of applications. This community-driven approach not only accelerates the development of VERSA but also ensures that it reflects the collective needs and insights of the global sound generation research community. Through active collaboration, VERSA aims to establish itself as a dynamic and enduring resource for the advancement of sound evaluation technologies.
@@ C.3 Speech Enhancement Evaluation Demonstration
The evaluation of speech enhancement is demonstrated in Table 7. While this paper focuses on six specific metrics, we anticipate broader adoption of the task by leveraging the diverse set of 65 metrics available in VERSA.